OpenStudy (anonymous):

assume that "diagonals are not congruent in a rectangle". By definition of rectangle, it has four right angles and opposite sides are congruent. So, by pythagorean theorem, AB^2 + BC^2 = AC^2 and BC^2 + CD^2 = BD^2. Since opposite sides are congruent, AB=CD, AD = BC. By substitution BC^2 = BD^2. that means diagonals are congruent. Thats a contradiction to what we have assumed. so the initial assumption must be wrong, and the opposite of that must be correct. So, the diagonals of rectangle are congruent. @ganeshie8

OpenStudy (anonymous):

can you help with this please

ganeshie8 (ganeshie8):

yeah ?

OpenStudy (anonymous):

sorry i copied the wrong thing

ganeshie8 (ganeshie8):

:) i guessed that lol

OpenStudy (anonymous):

Matt is constructing two similar triangles for an art project. The drawing below shows Matt's plans, but there is an error in his drawing. What changes would he make to the dimensions to change the error? Explain your reasoning using complete sentences.

ganeshie8 (ganeshie8):

find the ratios of corresponding sides first

ganeshie8 (ganeshie8):

if the triangles are similar we should get same number for all ratios

OpenStudy (anonymous):

ok how wwould i go about finding the ratios

ganeshie8 (ganeshie8):

we should get below ratios same : \(\large \frac{13}{2.5} = \frac{20.8}{4} = \frac{43}{8.4}\)

ganeshie8 (ganeshie8):

use ur calculator and simplify each ratio

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\large \frac{13}{2.5} = \frac{20.8}{4} = \frac{43}{8.4}\) \(\large 5.2 = 5.2 = 5.119\)

ganeshie8 (ganeshie8):

clearly, first two ratios are matching exactly. there seems to be a problem wid third ratio, eh ?

OpenStudy (anonymous):

yea thats wat i got

ganeshie8 (ganeshie8):

since the correct ratio is \(5.2\), we want third ratio also to equal this : say, we want to change the side DF from 43 to fix this error, then : \(\large \frac{DF}{8.4} = 5.2\) \(\large DF = 43.68\)

ganeshie8 (ganeshie8):

So, to fix the error, we need to change DF from 43 to 43.68

OpenStudy (anonymous):

ok and how can we do that?

ganeshie8 (ganeshie8):

we're done !

ganeshie8 (ganeshie8):

the question is just asking us to find out wat changes he needs to do to the dimensions.

ganeshie8 (ganeshie8):

we concluded he needs to change DF from 43 to 43.68. so we're done !

OpenStudy (anonymous):

ooh ok well thank you but i have more if you have the time

ganeshie8 (ganeshie8):

post them in another thread il try if im awake... it 1am here in india lol :)

OpenStudy (anonymous):

ooh